OpenStudy (anonymous):

Write y = 0.2x - 0.3 in standard form using integers. Select one: a. 2x + 10y = -3 b. -2x + 10y = -3 c. -2x - 10y = -3 d. 2x - 10y = -3

OpenStudy (kc_kennylau):

what do you get after subtracting 0.2x from both sides? :)

OpenStudy (anonymous):

.1 ?

OpenStudy (anonymous):

i was lost on this one

OpenStudy (kc_kennylau):

eh... i mean from both sides of y=0.2x-0.3

OpenStudy (anonymous):

i had D for the answer was that right ?

OpenStudy (kc_kennylau):

no :/

OpenStudy (kc_kennylau):

solve it step by step, as one of the teachers from my school says: "Step by step, no regret" :D

OpenStudy (anonymous):

dang , can you give me the answer and steps so i can see on the next questions ? Its a prep test HW

OpenStudy (kc_kennylau):

ok :)

OpenStudy (kc_kennylau):

So we have:\[y=0.2x-0.3\] Subtracting \(0.2x\) from both sides and we get:\[y-0.2x=0.3\] Multiplying \(10\) to both sides and we get:\[10y-2x=3\] Changing the terms on the left hand side and we get:\[-2x+10y=3\] P.S. \(10y-2x\) is the same as \(+10y-2x\) :D

OpenStudy (kc_kennylau):

hope that helps you :D

OpenStudy (kc_kennylau):

soorry it's -0.3 and -3 instead of 0.3 and 3

OpenStudy (kc_kennylau):

so the answer is b

OpenStudy (anonymous):

otayy , thanks

OpenStudy (kc_kennylau):

no problem :)